Student tuition at Boehring University is $150 per semester credit hour. The s
tate supplements school revenue by $100 per semester credit hour. Average
class size for a typical 3-credit course is 50 students. Labor costs are $4,000 p
er class, material costs are $20 per student per class, and overhead costs are
$25,000 per class.
a. What is the multifactor productivity ratio for this course process?
b. If instructors work an average of 14 hours per week for 16 weeks for each
3-credit class of 50 students, what is the labor productivity ratio?
1
Expert's answer
2020-10-19T08:33:59-0400

a) Multifactor productivity = Value of Output/ Value of input resources

 Value of Output = 50 × 3 × ($150 + $100) = 50 × 3 × $250 = $37,500 /class

 Value of Inputs = Labor cost + Material cost + Overhead cost = $4000 + ($20 × 50) + $25,000 = $30,000 /class

Therefore, the multifactor productivity = $37,500 / $30,000 = 1.25